3. Consider a system with a following transfer function: G(s) %3D s2 + 2s + 1 a. What are the state equations of the system iPlease show all work, really need to double check this one.
Answer #1

Б.Сc34 ] @ 44 - л а, ла, In controlisle canonical kolm, <a = 2 * (*) 8= CG] Хан - ( + Pч ч с (+ O2x 3) 4-ки к: Cқ к] Е, Е. |al $+(2+62) 5+ (1+K,)=0 @ for contically damped systems, 87 246 Wh stwin = o. 2 i + 2 what win = 0 (sthens = o comparing equa